Name: Grading “Rubric” – The paper:

1. (25) Clearly distinguishes between “They say” and “I say” sections

2. (25) Accurately presents facts and viewpoints of the sources

3. (25) Has a clearly articulated response or personal viewpoint in the ”I say” section 4. (10) Is clear and direct in its use of language

5. (5) Avoids run-on or overly complicated sentences

6. (5) Is free of jargon, cliches, and misused words

7. (5) Avoids misspellings, grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns), uses correct punctuation

Total: